Suppose a perfect competitive market comprise 1000 firms.
If the market demand and supply functions are following;
Qd= 170,000,000 - 10,000,000p
Qs= 70,000,000 + 15,000,000.
a) calculate the Pe and We
b) Given your answer in a, how much output will be produced by each firm in the industry.
c) suppose that one of the firm goes out, what will be the effect on Pe and Qe.
1
Expert's answer
2021-03-31T11:46:14-0400

a)


"170,000,000-10,000,000p=70,000,000+15,000,000p"

"p=4"

"Q=130,000,000"

b)


"\frac {130,000,000}{1,000}=130,000"

c)


"\frac {130,000,000}{999}=130,130"